Find the distance between (-9,2) and (-3,-6)
Answer by JP(22)

First, since you are trying to find the distance between those points you must use something called the "distance formula". The distance formula is derived from the pythagorean theorem. But to the point, the distance formula is d=sqrt(x2-x1)^2+(y2-y1)^2
You have the points (-9,2)(-3,-6). So: -9=x1, 2=y1, -3=x2, -6=y2 --now plug them in
d=sqrt(-3--9)^2+(-6-2)^2 now be careful when doing the operations as to not get caught up by simple mistakes.
-3--9 is like saying -3++9=6
-6-2 is like saying -6+-2=-8
You should now get...d=sqrt(6)^2+(-8)^2
6^2=36
-8^2=64 because when you multiply two negatives together they form a positive number. -8*-8=64
So now you should have...d=sqrt36+64
Now you simply add what's left to get 36+64=100
Now you have d=sqrt100
The square root of 100 equals 10
so d=10
To check your solution simply plug in 10 for d
so you have 10=sqrt(-3--9)^2+(-6-2)^2
and your solution will be...
10=10
Hope this helps...